# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

### v1

@@ 5.3 PEP Hard Idiom Benchmark Performance
Untrained 0.538 0.2637 0.3539 0.1396 0.1479 0.1436
SFT 0.7686 0.3178 0.4497 0.2976 0.296 0.2968
SFT+RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
CoT Untrained 0.8812 0.6854 0.771 0.5049 0.4878 0.4962
CoT (RS-SFT) 0.935 0.8183 0.8727 0.6639 0.65 0.6569
CoT (RS-SFT + RS-DPO) 0.9234 0.8643 0.8929 0.771 0.7571 0.764
Table 1: MetaLint Qwen3-4B results across training conditions evaluated on the Ruff meta-linting task.
Model In-Domain Near Transfer Far Transfer
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
Untrained 0.45 0.14 0.22 0.58 0.24 0.34 0.54 0.29 0.38
SFT 0.93 (+0.48) 0.74 (+0.6) 0.83 (+0.61) 0.89 (+0.31) 0.24 (+0) 0.38 (+0.04) 0.72 (+0.18) 0.27 (-0.02) 0.39 (+0.01)
SFT+RS-DPO 0.72 (+0.27) 1 (+0.86) 0.83 (+0.61) 0.76 (+0.18) 0.8 (+0.56) 0.78 (+0.44) 0.75 (+0.21) 0.81 (+0.52) 0.78 (+0.4)
CoT Untrained 0.87 0.5 0.63 0.95 0.88 0.91 0.87 0.68 0.76
CoT (RS-SFT) 0.87 (+0) 0.73 (+0.23) 0.8 (+0.17) 0.97 (+0.02) 0.86 (-0.02) 0.91 (+0) 0.94 (+0.07) 0.82 (+0.14) 0.88 (+0.12)
CoT (RS-SFT + RS-DPO) 0.86 (-0.1) 0.85 (+0.35) 0.85 (+0.22) 0.97 (+0.02) 0.92 (+0.04) 0.94 (+0.03) 0.92 (+0.05) 0.86 (+0.18) 0.89 (+0.13)
Table 2: MetaLint Qwen3-4B detection results across training conditions evaluated on the Ruff meta-linting task across transfer conditions (In-Domain, Near Transfer and Far Transfer)
@@
Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.354
Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.163 0.1477 0.155
@@
GPT-4o 0.8938 0.6788 0.7716 0.4461 0.332 0.3807
GPT-4.1 0.907 0.646 0.7546 0.4632 0.4673 0.4653
@@
6 Conclusion
Our results demonstrate that MetaLint may encourage more adaptive reasoning over idiom specifications instead of rote memorization of code patterns through instruction-following and easy-to-hard generalization from synthetic data. By training on synthetic examples of easy-to-detect idioms derived from linter signals, MetaLint enables both Qwen3-4B CoT and non-CoT variants to generalize well across Ruff idioms in in-domain, near-transfer, and far-transfer settings, as well as to real-world, hard-to-detect PEP idioms. The models show substantial improvements in both detection and localization, particularly after RS-DPO training, with the non-CoT variant achieving the highest detection recall among all evaluated models, despite its smaller size and training solely on synthetic data. Additionally, it achieves this with just 4B parameters and no test-time compute in the non-CoT setting, making it one of the most efficient models. These findings underscore the effectiveness of synthetic instruction-tuned training for fostering generalization and reasoning, even in low-resource scenarios where annotated examples of harder idiom categories are limited or unavailable. Finally, we note that for easy idioms that are more “mechanical” to detect, linters remain the most cost-effective option. However, LLMs—when trained with the MetaLint approach—enable detection of more abstract idioms, potentially unlocking support for personalized linting, and offering a no-code solution for code quality enforcement.
7 Limitations and Future Work
Despite the promising results achieved by MetaLint, our work has several limitations that we plan to address in future research. First, our experiments are restricted to a single model family (Qwen3) and model size (4B). We intend to extend our evaluations to other model families, including LLaMA and Phi, as well as larger Qwen models (8B and 14B). In the CoT setting, it remains an open question whether non-CoT models can be trained to effectively produce CoT-style reasoning with supervision from a teacher model. We also explored self-improvement strategies for RS-SFT data generation in cases where the base model failed, such as STaR (Zelikman et al., 2022), but found it challenging to generate CoTs that do not directly reference provided hints, which risks contaminating the training data. As a result, we adopted a simpler rejection sampling strategy. Furthermore, our approach does not yet incorporate more advanced reinforcement learning techniques such as direct GRPO (Shao et al., 2024) using our verifiable linter-based reward model, or curriculum learning methods to control the progression of idiom difficulty within synthetic training data. Our current experiments also focus exclusively on the Python programming language. Future work will explore extensions to other languages such as Java using Java-specific linters, as well as cross-language generalization by training on Python idioms and evaluating on Java idioms, and vice versa. Finally, while we prompt and train the models to generate refactoring fixes for code with idiom violations to improve code quality, we do not evaluate this capability in the present work and leave it for future exploration.
Author Contributions
Atharva Naik: Conception, Ruff Idiom Data Generation, SFT, RS-SFT, RS-DPO Experiments, Baselines, Writing
Lawanya Baghel: RS-DPO Experiments, PEP Hard Idiom Benchmark Creation, Writing
Dhakshin Govindarajan: PEP Hard Idiom Benchmark Creation
Darsh Agrawal: PEP Hard Idiom Benchmark Creation
Daniel Fried: Advising, Conception, Writing
Carolyn Rose: Advising, Conception, Writing
@@ Acknowledgments
Appendix A Appendix
A.1 Method Additional Details
A.1.1 MetaLint Instruction Following Prompt
@@
Example Input Look at the following list of code idiom specifications with definitions and examples: # Idiom ANN202 (missing-return-type-private-function) Definition: Checks that private functions and methods have return type annotations. Rationale: Type annotations are a good way to document the return types of functions. They also help catch bugs, when used alongside a type checker, by ensuring that the types of any returned values, and the types expected by callers, match expectation. Example:
@@
Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say ’NO VIOLATIONS FOUND’ if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file:
@@
Example Output **Idiom ANN202 Violations:**
@@
A.1.2 DPO Contrastive Pair and RS-SFT Sampling Details
@@
For RS-DPO sampling (in both CoT and non-CoT settings), we used the standard MetaLint instruction-following prompt with the SFT models. In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded “Baseline Inference Prompt” described in Section A.1.4.
A.1.3 Training Hyperparameters and Computational Environment
We use the following hyperparameters for the SFT and DPO training experiments.
SFT/RS-SFT hyperparameters:
For both SFT and RS-SFT we fine-tune the Qwen3-4B model using flash_attention_2 and bfloat16 precision. The model is trained for 2 epochs with a learning rate of 2e-5, cosine learning rate schedule, and a warmup ratio of 0.1. We use a maximum sequence length of 3000 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing is enabled to reduce memory usage, with non-reentrant mode. Evaluation is performed every 2000 steps, and checkpoints are saved at the same interval. Special tokens are manually handled in the chat template without automatic insertion. The training uses 12 preprocessing workers and is seeded with 42 for reproducibility.
RS-DPO parameters:
@@ A.1.3 Training Hyperparameters and Computational Environment
Computational Environment: All SFT, RS-SFT and RS-DPO experiments were conducted on a Linux server equipped with two NVIDIA A100 80GB PCIe GPUs (Ampere architecture), CUDA 12.9, and driver version 575.51.03. Each job had access to 100 GB of CPU memory and 2 CPU cores. Training was performed using mixed-precision (bfloat16) with gradient checkpointing to optimize memory usage. For inference we used a similar environment but varied number of GPUs based on model sizes and needs.
A.1.4 Baseline Inference Details
@@
Closed Source LLMs: We use the chat completion OpenAI API with max tokens of 1024 for GPT-4.1 and GPT-4o and max completion tokens of 3000 for o3-mini and o4-mini. We use default parameters for everything else (temperature of 1 and top-p of 1, no presence penalty).
@@
Baseline Inference Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file: {CODE_FILE} # OUTPUT FORMAT I want you to generate your output under a section called “### Final Idiom Violations Found”. Structure you response for a given idiom XYZ as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:**
@@
Inquiry.objects.filter(owner=kwargs.pop(’user’)))", "fix": null}

### v2

@@ 5.1 Generalization on Synthetic Data
Qwen3-4B 0.5380 0.2637 0.3539 0.1396 0.1479 0.1436
Qwen3-4B + SFT 0.7686 0.3178 0.4497 0.2976 0.2960 0.2968
Qwen3-4B + SFT + RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
Qwen3-4B w CoT 0.8812 0.6854 0.7710 0.5049 0.4878 0.4962
Qwen3-4B w CoT + RS-SFT 0.9350 0.8183 0.8727 0.6639 0.6500 0.6569
Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9234 0.8643 0.8929 0.7710 0.7571 0.7640
Table 1: Cross-Idiom Generalization on Python Ruff Idioms: Effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning). Best score across the compared training setups per model are bolded.
PMD and JEP Tree-Sitter Idioms: To demonstrate the generality of MetaLint training across programming languages and linters, we present results from training on PMD and JEP Tree-Sitter synthetic data in Table 2 (full results in Table 28). Training on PMD shows the same overall pattern as before but with larger recall gains for both SFT and DPO, and notably stronger localization under DPO. For Llama3.1‑8B‑Instruct, SFT initially reduces detection precision, which DPO then recovers; the same precision dip‑and‑recovery appears when transferring PMD→JEP for Llama3.2‑3B‑Instruct. Despite never seeing JEP idioms during training, DPO models achieve strong detection and localization on JEP. In the untrained setting, Llama3.2‑3B‑Instruct (on PMD) and Llama3.1‑3B‑Instruct (on JEP) nearly always output the correct format but predict NO VIOLATIONS FOUND, yielding zero or near‑zero scores because our metrics exclude that class for detection and only score positive cases for localization. Training on JEP yields high in‑domain performance for all metrics with minimal additional benefit from DPO, likely due to JEP’s smaller idiom set (15 vs 269 for PMD) and more precise instructions (Table 5). In the harder JEP→PMD transfer, DPO outperforms SFT, though overall transfer remains weaker than PMD→JEP, reflecting PMD’s broader diversity and more challenging specifications (Appendix C.5).
Overall, MetaLint training consistently yields more adaptable models than the base model, but performance depends on the diversity of training idioms and the gap in instruction quality between training and test data.
Model Transfer Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
Llama3.2-3B-Instruct PMD $\rightarrow$ PMD 0.0457 0.0079 0.0134 0.0015 0.0022 0.0017
Llama3.2-3B-Instruct + SFT 0.2251 0.4421 0.2983 0.2822 0.2778 0.2800
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4395 0.8908 0.5886 0.5930 0.5969 0.5949
Llama3.2-3B-Instruct PMD $\rightarrow$ JEP 0.3855 0.0096 0.0187 0.0005 0.0004 0.0005
Llama3.2-3B-Instruct + SFT 0.2286 0.4072 0.2928 0.1626 0.1336 0.1467
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4903 0.8338 0.6175 0.4216 0.3333 0.3721
Table 2: Cross-Idiom Generalization on JEP & PMD Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 28). The transfer column indicates training and test data on the left and right side of the arrow. Best score across the compared training setups per model are bolded.
5.2 Evaluating Easy-to-Hard Generalizations
To evaluate whether MetaLint training on easy, linter-detectable Ruff idioms improves performance on hard, manually curated PEP idioms, we report results on our PEP hard idiom benchmark (Table 3, full results in Table 19). At the SFT stage, performance declines for Qwen3-4B (with and without CoT) but improves slightly for Llama3.2-3B-Instruct, suggesting that SFT can induce memorization of the training distribution and reduce adaptability. In contrast, DPO yields clear improvements in detection and localization (except detection precision for Llama3.2-3B-Instruct), with statistically significant gains (Appendix E.2). An additional experiment training Qwen3-4B (CoT) directly with RS-DPO, bypassing SFT, resulted in near-zero performance because many generated DPO pairs violated the required output format, which the model inherited. Thus, SFT, despite its drawbacks, is essential for teaching format compliance and setting the stage for DPO to unlock easy-to-hard generalization. Interestingly, the non-CoT model achieves substantially higher detection recall and slightly higher F-score than the CoT variant, despite lower precision. Our analysis attributes the CoT model’s reduced recall to its more conservative interpretation of idiom specifications and to errors such as misinterpretation, overthinking, and skipped lines, as detailed in Appendix E.3.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B + SFT 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
Qwen3-4B + SFT + RS-DPO 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
Qwen3-4B w CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-4B w CoT + RS-SFT 0.7615 0.3689 0.4970 0.2785 0.1437 0.1896
Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
Table 3: Easy-to-Hard Generalization on PEP Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3). Best score across the compared training setups per model are bolded.
5.3 Benchmarking on Hard Idioms
Table 4 compares the best-performing Qwen3-4B MetaLint DPO models against state-of-the-art code and reasoning models (full results in Table 17).
Detection: In terms of detection F-score, the non-CoT MetaLint model is competitive with o3-mini and GPT-5 but is outperformed by some larger open-source models (e.g., Qwen3-32B with CoT, DeepSeek-R1-Distill-Qwen-32B with CoT, and GPT-oss-120B) and closed-source models (GPT-4o, GPT-4.1, and o4-mini). However, the non-CoT model achieves the highest detection recall among all evaluated models, while the CoT model ranks among the top in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
Localization: For localization, the MetaLint models lag behind larger 32B and 120B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix E.2) and outperform GPT-oss-20B. This is notable given that the MetaLint models are much smaller (4B parameters), trained only on synthetic data derived from easy idioms, and that the non-CoT model does not use test-time compute.
Overall, the strong results, especially the best-in-class recall of the non-CoT model, demonstrate the effectiveness of our framework in achieving easy-to-hard generalization. This is enabled by training on synthetic data with easy idioms and by encouraging adaptive reasoning through instruction fine-tuning and DPO rather than relying on rote memorization.
@@ 5.3 Benchmarking on Hard Idioms
Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.3540
Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.1630 0.1477 0.1550
@@
DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.7130 0.4015 0.3403 0.3684
GPT-oss-20b 0.8377 0.3531 0.4968 0.2510 0.1695 0.2024
GPT-oss-120b 0.9157 0.6456 0.7573 0.3991 0.3331 0.3631
Qwen3-4B MetaLint (SFT+RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
@@
GPT-4o 0.8938 0.6788 0.7716 0.4461 0.3320 0.3807
GPT-4.1 0.9070 0.6460 0.7546 0.4632 0.4673 0.4653
GPT-5 (high) 0.9130 0.5673 0.6998 0.4397 0.4257 0.4326
Table 4: Benchmarking on Hard Idioms: Results comparing state of the art code and reasoning models on the hard PEP benchmark to contextualize the gains achieved with MetaLint training. The best scores are bolded and second best and underlined.
6 Conclusion and Future Work
Our results show that MetaLint training fosters adaptive reasoning over idiom specifications rather than rote memorization. We observe generalization to unseen idioms in Python and Java, across three linters (Ruff, PMD, JEP tree-sitter), two model families (Qwen, Llama), reasoning and non-reasoning settings, and multiple scales (3B, 4B, 8B). Easy-to-hard generalization occurs from linter-detectable Ruff idioms to harder PEP idioms, with SFT teaching output formatting and DPO enabling true generalization. Compared to state-of-the-art code and reasoning models, MetaLint-trained Qwen models have detection comparable with o3-mini and GPT-5, achieving highest recall (non-CoT) and third-best precision (CoT). Localization lags but surpasses GPT-oss-20B with only 4B parameters and no test-time compute and is comparable to o3-mini, demonstrating efficiency. These results highlight the effectiveness of instruction fine-tuning and preference optimization on synthetic data for reasoning and generalization, even with scarce annotated examples. For mechanically easy idioms, linters remain cost-effective, but MetaLint enables detection of abstract idioms, supporting personalized, evolving code quality standards. We plan to release code and data for reproducibility. Future work includes training for automated refactoring and exploring advanced RL methods like Group Relative Policy Optimization (GRPO) Shao et al. (2024).
@@ Acknowledgments
Appendix A Limitations
Despite the promising results achieved by MetaLint, our work has some limitations that we plan to address in future research. For the CoT setting, we didn't explore whether non-CoT models can be trained to effectively produce CoT-style reasoning with supervision from a teacher model. We also explored self-improvement strategies for RS-SFT data generation in cases where the base model failed, such as STaR (Zelikman et al., 2022), but found it challenging to generate CoTs that do not directly reference provided hints, which risks contaminating the training data. As a result, we adopted a simpler rejection sampling or RS-SFT strategy. Furthermore, our approach does not yet incorporate more advanced reinforcement learning techniques such as Group Relative Policy Optimization (GRPO) (Shao et al., 2024) using our verifiable linter-based reward model, or curriculum learning methods to control the progression of idiom difficulty within synthetic training data. Our current experiments also focus on training on one language at a time, such as only Python or Java. Future work will explore joint training and extension to more programming languages like JavaScript, Ruby, Go, etc., as well as cross-language generalization by training on Python idioms and evaluating on Java idioms, and vice versa. Finally, while we do not evaluate or train for refactoring of the idiom-violating code, we plan to do so in future work.
Appendix B More Related Work
Easy-to-Hard Generalization. Research shows that training on simpler problems enhances generalization to harder ones in math, algorithms, and code, motivating its application to code quality analysis. In math reasoning, models trained on easier problems (e.g., level 1–3) consistently generalize better to harder benchmarks (e.g., level 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025). Several works emphasize the importance of selecting high-quality supervision for harder problems (He et al., 2024). Beyond math, Sun et al. (2024a) shows that reward models trained on simple code and math problems improve performance on complex ones. Broader studies on multi-task and length generalization (Hu et al., 2025) and differentiable programming (Gaunt et al., 2016) reveal how structural simplicity during training can lead to robustness on longer or more complex reasoning instances, including code. Zhang et al. (2024a) reinforces this by evaluating reward models on algorithmic tasks like string manipulation and demonstrating transfer from simpler to harder formats. Drawing inspiration from this work, we train MetaLint on large-scale synthetic data covering easily detectable code idioms handled by rule-based linters, and hypothesize that these simple patterns serve as stepping stones toward generalizing to complex, novel PEP idioms.
Appendix C Method Additional Details
C.1 MetaLint Instruction Following Prompt
@@
Example Ruff Meta-Task Input Look at the following list of code idiom specifications with definitions and examples: # Idiom ANN202 (missing-return-type-private-function) Definition: Checks that private functions and methods have return type annotations. Rationale: Type annotations are a good way to document the return types of functions. They also help catch bugs, when used alongside a type checker, by ensuring that the types of any returned values, and the types expected by callers, match expectation. Example:
@@
Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say 'NO VIOLATIONS FOUND' if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file:
@@
Example Ruff Meta-Task Output **Idiom ANN202 Violations:**
@@
C.2 DPO Contrastive Pair and RS-SFT Sampling Details
@@
For RS-DPO sampling (in both CoT and non-CoT settings), we used the standard MetaLint instruction-following prompt with the SFT models. In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded ``Baseline Inference Prompt'' described in Section C.4.
C.3 Training Hyperparameters and Computational Environment
Python SFT/RS-SFT hyperparameters:
We fine-tune the Qwen3-4B model using flash_attention_2 and bfloat16 precision. The model is trained for 2 epochs with a learning rate of 2e-5, cosine learning rate schedule, and a warmup ratio of 0.1. We use a maximum sequence length of 3000 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing is enabled to reduce memory usage, with non-reentrant mode. Evaluation is performed every 2000 steps, and checkpoints are saved at the same interval. Special tokens are manually handled in the chat template without automatic insertion. The training uses 12 preprocessing workers and is seeded with 42 for reproducibility.
Python RS-DPO parameters:
@@ C.3 Training Hyperparameters and Computational Environment
Java SFT hyperparameters:
For Java experiments, we fine-tune Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct with bfloat16 precision. Both models are trained for 2 epochs with a learning rate of 2e-5, cosine learning rate schedule, and warmup ratio of 0.1. We use a maximum sequence length of 3000 tokens, per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing (non-reentrant) is enabled. Evaluation and checkpoint saving occur every 5000 steps. Special tokens are manually handled in the chat template. Training is seeded with 42.
Java RS-DPO parameters:
RS-DPO training is performed on Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct using bfloat16 precision. Training runs for 1 epoch with a learning rate of 5e-7, cosine learning rate scheduling, and warmup ratio of 0.1. We use a maximum input length of 3500 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing (non-reentrant) is enabled. Evaluation and checkpoints are recorded every 200 steps. Reward shaping parameters vary across settings, with $\beta\in\{0.1,0.5,1\}$. Seeds are fixed at 42 for reproducibility.
Computational Environment:
All SFT, RS-SFT, and RS-DPO experiments (Python and Java) were conducted on a Linux server equipped with NVIDIA A100 80GB GPUs (Ampere architecture), CUDA 12.9, and driver version 575.51.03. Each job had access to 100 GB of CPU memory and 2 CPU cores. Training used mixed-precision (bfloat16) with gradient checkpointing to optimize memory usage. Inference used a similar setup with GPU allocation varying by model size.
C.4 Baseline Inference Details
@@
Closed Source LLMs: We use the chat completion OpenAI API with max tokens of 1024 for GPT-4.1 and GPT-4o and max completion tokens of 3000 for o3-mini and o4-mini. We use default parameters for everything else (temperature of 1 and top-p of 1, no presence penalty). For GPT-5 we use 8192 max completion tokens and high reasoning effort.
@@
Baseline Inference Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file: {CODE_FILE} # OUTPUT FORMAT I want you to generate your output under a section called ``### Final Idiom Violations Found''. Structure you response for a given idiom XYZ as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:**
@@
Inquiry.objects.filter(owner=kwargs.pop('user')))", "fix": null}